Exercise 8-14 The following is a record of Pervis Ellison Companys transactions for Boston Teapots for the month of May 2017 May 1 Balance 400 units @ $20 May 10 Sale 300 units $38 12 Purchase 600 units $25 28 Purchase 400 units $30 20 Sale 540 units$38 Your answer is incorrect. Try again. Assuming that perpetual inventories are not maintained and that a physical count at the end of the month shows 560 units on hand, what is the cost of the ending inventory using (1) FIFO and (2) LIFO? FIFO LIFO Ending Inventory
Answer #1
Periodic:
Ending Inventory:
FIFO 16000 =(400*30)+(160*25)
LIFO 12000 =(400*20)+(160*25)
Perpetual:
Ending Inventory:
FIFO 16000
LIFO 15500 =(400-300)*20+(600-540)*25+(400*30)